OpenStudy (anonymous):

Two triangles can be formed with the given information. Use the Law of Sines to solve the triangles. B = 40°, b = 25, c = 26

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

yep... 1 triangle will have an angle C as an acute angle, an the other will have angle C as an obtuse angle.

OpenStudy (anonymous):

Okay this one is tough... I need the values of A C and a on the first triangle, and A C and a on the second one

OpenStudy (jdoe0001):

get one first

OpenStudy (anonymous):

Idk what to do, like, at all

OpenStudy (jdoe0001):

well, you do know the Law of Sines, right?

OpenStudy (anonymous):

a/sinA = b/sinB = c/sinC

OpenStudy (jdoe0001):

right, so use the "b" and "B" values and "c" given using the part b/sinB = c/sinC

OpenStudy (jdoe0001):

sinB/b = sinC/c either way, you can flip them around btw

OpenStudy (anonymous):

a/sinA = 25/sin40 = 26/sinC

OpenStudy (jdoe0001):

yes, the idea is that you may not have all 3 at once, just a pair, so you use just one pair at a time :)

OpenStudy (jdoe0001):

so just use 25/sin40 = 26/sinC

OpenStudy (jdoe0001):

so, what do you get for the sinC?

OpenStudy (anonymous):

Im trying but idk how to do it

OpenStudy (jdoe0001):

well, at 25/b = 26/c how would you solve for "c"?

OpenStudy (anonymous):

move the left to the right?

OpenStudy (campbell_st):

here is the 1st case \[\frac{Sin(40)}{25} = \frac{\sin(C)}{26} \] solve for C C = 41.95 degrees and you also need to consider the case when C = 180 - 41.95 so C = 138.05 degrees. 

OpenStudy (campbell_st):

find angle A by angle sum of a triangle when you know the values of angle C

OpenStudy (anonymous):

A = 98.05?

OpenStudy (anonymous):

the other A = 1.95?

OpenStudy (campbell_st):

well that seems to work.

OpenStudy (anonymous):

but then how do i get a?

OpenStudy (campbell_st):

well if you need a then \[\frac{25}{\sin(40)} = \frac{a}{\sin(98.05)}..... and ..... \frac{25}{\sin(40)} = \frac{a}{\sin(1.95)}\] solve for a. each triangle will have a different value for a. in the 1st case a will be larger than 26 in the 2nd triangle a will be very small... hope this helps.

OpenStudy (anonymous):

It helps a whole lot man. Thank u so much camp!

OpenStudy (anonymous):

A = 98°, C = 42°, a = 38.5; A = 2°, C = 138°, a = 1.4